# A Nonparametric Bayesian Local-Global Model for Enhanced Adverse Event Signal Detection in Spontaneous Reporting System Data

### v2

@@ 7 DISCUSSION
There are several limitations with our current approach that motivate interesting future directions. First, we focus here primarily on signal detection, i.e., to identify critical AEs of concern for each drug. The clusters of AEs produced by our DP priors have not been investigated but may reveal additional, deeper insights by identifying groups of “similar AEs”. A reformulation via mixed membership models could further enhance interpretability by allowing the same AE to appear in multiple clusters/groups. Second, while we consider data-driven characterization of similarities in AE signal strengths between drugs, a more flexible approach acknowledging known Drug-Drug Interactions (DDI poleksic2019database ), possibly via appropriate prior distributions, may further enhance signal detection. Finally, our current framework aggregates drug-AE data across time points for a cross-sectional analysis. Extending the model to accommodate temporal variations in SRS data could enable formal identification of evolving AE patterns.
@@
$\lambda_{ij}\sim z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g}$
@@
$\lambda_{ij}\sim z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g}$
@@
S.1.2 Some notes on the “global-only” model
A special case of our local-global models leads to a reduced “global-only” model when $\pi^{g}=0$, i.e., $\lambda_{ij}=\lambda_{i}^{g}$ for $j=1,\dots,J-1$. Under this model there is no provision for different $\lambda_{ij}$ for any $j\neq J$. The global model has a simplified likelihood $n_{ij}\sim\text{Poisson}(\lambda_{i}E_{ij})$ here $\lambda_{i}^{g}\sim D$ if $j\neq J$, which produces $\sum_{j=1}^{J-1}n_{ij}\sim\text{Poisson}(\lambda_{i}\sum_{j=1}^{J-1}E_{ij})$. The model reduction also impacts the signal detection. For the local-global framework, hypothesis testing is conducted for each individual $\lambda_{ij}$. On the other hand, under the global model, the test is simplified to evaluating whether $\lambda_{i}^{g}>1$. Testing $\lambda_{i}^{g}>1$ entails assessing if the summarized Poisson count $\sum_{i}n_{ij}$ is large compared to the expected count $\sum_{i}E_{ij}$, owing to the reduced model. This approach represents a “global test”, whereas our proposed local-global model allows conducting hypothesis testing for each individual drug-AE pair. As a result, directly comparing the global model with the local-global model is not feasible, as they target different types of assessment.
@@
In this section, we first provide MCMC algorithms implementing our models in Section S.2.1. The MCMC algorithms are derived using finite mixture approximation of the underlying DP processes [23, 16], and combine Gibbs sampling updates with stepping-out slice sampling updates [31] for various parameters. Some additional notes on the MCMC implementation, including initial value generation for our MCMC algorithms are then provided in Section S.2.2. Next we provide detailed notes on the grid-search method identifying optimal quantile-level $p$ and threshold $k$ used for our hypothesis tests for signal detection (Section S.2.3). The JAGS code [32] for Hu et al.’s model [18] is shown as a reference in Section S.2.4. The example statin dataset is included in Section S.2.5.
@@
    Here $f_{\text{Gamma}}(\theta\mid\text{shape}=\beta,\text{rate}=\beta)$ denotes the probability density function of the Gamma($\beta,\beta$) distribution at $\theta$. $\pi(\beta)$ is the prior distribution of $\beta$. Direct random sampling from this density is not feasible. Instead, we use stepping-out slice sampling [31] to generate Markov chain draws for $\beta$.
@@
Algorithm S.2 One Step Iteration of DP Poisson local-global column model (Given data $\{(n_{ij},E_{ij},z_{ij}^{l}):i=1,\dots,I;j=1,\dots,J\}$ )
@@
    Generate global $\text{DPSB}_{K}(\bm{\eta},\bm{\theta},\alpha,\beta)$ parameters using Algorithm S.1, with $\{(\tilde{n}_{i},\tilde{E}_{i},1(\sum_{j=1}^{J-1}z_{ij}^{l}<J-1)):i=1,\dots,I\}$ as data where
    $\tilde{n}_{j}=\sum_{j=1}^{J-1}n_{ij}(1-z_{ij}^{l}),\quad\tilde{E}_{j}=\sum_{j=1}^{J-1}E_{ij}(1-z_{ij}^{l}).$
@@
    $z_{ij}^{l}\mid n_{ij},E_{ij},z_{ij}^{l},\lambda_{ij}^{l},\lambda_{i}^{g}\sim\text{Bernoulli}(p_{ij}^{l}),$
@@
    $p_{ij}^{l}=\frac{\pi^{l}f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})}{\pi^{l}f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})+(1-\pi^{l})f_{\text{Poisson}}(n_{ij}\mid\lambda_{i}^{g}E_{ij})}.$
@@
    Define $\lambda_{ij}=z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g}$ for all $i=1,\dots,I$ and $j=1,\dots,(J-1)$.
@@
    Generate $\pi^{l}$ using:
    $\pi^{l}\mid z_{ij}^{l}\sim\text{Beta}\left(a^{l}+\sum_{i=1}^{I}\sum_{j=1}^{J-1}z_{ij},b^{l}+\sum_{i=1}^{I}\sum_{j=1}^{J-1}(1-z_{ij})\right),$
    Notes: $a^{l}$ = 1, $b^{l}$ = 1
@@
Algorithm S.3 One Iteration of DP ZIP local-global Model Given data $(n_{ij},E_{ij},z_{ij}^{l})$
@@
Given the observed data $n_{ij}$ and $E_{ij}$ for $i=1,\dots,I$, $j=1,\dots,J$, and a prespecified number $K$ of DP mixture components for finite approximation. The default value for $K$ in our software is set to $\max(\log(IJ),10)$, a setting that is generally sufficient for most cases. For the simulation study in Section 5, we used $K=10$, which satisfies the condition $K>\log(1492\times 7)$. In the data analysis presented in Section 6, we chose a more conservative value of $K=20$ to allow for a larger number of atoms in the DP, facilitating a more cautious interpretation of the results. We initialize the number of nonempty components $K_{+}$ to the smallest integer greater than $\min\{K/2,I-1\}$. The local-global indicators $z_{ij}^{l}$ are randomly initialized by independently sampling each cell $(i,j)$ from a $\text{Bernoulli}(0.5)$ distribution. The DP parameters for each local and global DP process are then initialized using the subset of data $(n_{ij},E_{ij})$ assigned to that process, as determined by the indicators $z_{ij}^{l}$, for $i=1,\dots,J-1$.
@@
Table S.1: The first 30 rows of the SRS contingency table summarized from the FDA FEARS database (Q3 2014-Q4 2020) which catalogs a total of 63,976,610 reported cases.
@@
We present detailed values of the simulation to compare among our drug (column)-based models (Poisson/ZIP), DP Hu et al., and posterior quantile optimized DP Hu et al. method. Four evaluation metrics (FDR, sensitivity, average type I error, and F-score) are evaluated as well as the main text with same simulation settings in Table 1. Frequentist and Bayesian MSE are compared to further evaluate the loss of the estimation.
@@
We present the four evaluation metrics in Tables S.2, S.3, S.4, and S.5 to provide a clearer numerical comparison beyond Figure 3. Additionally, we further explored the effect of posterior quantile optimization on the DP Hu et al. in the process of FDR control and minimizing FNR. This allows for a more comprehensive understand the strategy of optimizing the posterior quantile of the test statistic and deterministic threshold simultaneously.
@@
Case 1a
Weak correlation
No ZI
@@
3.0 0.0042 0.0042 0.0026 0.0056
Case 1b
Weak correlation
30% ZI
@@
3.0 0.0037 0.0037 0.0022 0.0049
Case 2a
Moderate correlation
No ZI
@@
3.0 0.0024 0.0023 0.0026 0.0056
Case 2b
Moderate correlation
30% ZI
@@
3.0 0.0025 0.0025 0.0029 0.0062
Case 3a
Strong correlation
No ZI
@@
3.0 0.0015 0.0017 0.0026 0.0058
Case 3b
Strong correlation
30% ZI
@@
Case 1a
Weak correlation
No ZI
@@
3.0 0.5245 0.5252 0.4818 0.4988
Case 1b
Weak correlation
30% ZI
@@
3.0 0.5256 0.5261 0.4813 0.4998
Case 2a
Moderate correlation
No ZI
@@
3.0 0.6597 0.6599 0.5667 0.5826
Case 2b
Moderate correlation
30% ZI
@@
3.0 0.6611 0.6613 0.5674 0.5835
Case 3a
Strong correlation
No ZI
@@
3.0 0.7959 0.7961 0.6307 0.6470
Case 3b
Strong correlation
30% ZI
@@
2.0 0.0000 0.0000 0.0000 0.0000
2.5 0.0000 0.0000 0.0000 0.0000
3.0 0.0000 0.0000 0.0000 0.0000
Case 2a
Moderate correlation
No ZI
@@
2.5 0.0000 0.0000 0.0000 0.0000
3.0 0.0000 0.0000 0.0000 0.0000
Case 2b
Moderate correlation
30% ZI
@@
Case 3a
Strong correlation
No ZI
@@
2.5 0.0000 0.0000 0.0000 0.0000
3.0 0.0000 0.0000 0.0000 0.0000
Case 3b
Strong correlation
30% ZI
@@
Case 1a
Weak correlation
No ZI
@@
3.0 0.6864 0.6869 0.6489 0.6635
Case 1b
Weak correlation
30% ZI
@@
3.0 0.6875 0.6879 0.6486 0.6646
Case 2a
Moderate correlation
No ZI
@@
3.0 0.7938 0.7939 0.7223 0.7342
Case 2b
Moderate correlation
30% ZI
@@
3.0 0.7948 0.7949 0.7227 0.7347
Case 3a
Strong correlation
No ZI
@@
3.0 0.8856 0.8856 0.7725 0.7836
Case 3b
Strong correlation
30% ZI
@@
3.0 0.8857 0.8858 0.7718 0.7829

### v3

@@ 7 DISCUSSION
There are several limitations with our current approach that motivate interesting future directions. First, we focus here primarily on signal detection, i.e., to identify critical AEs of concern for each drug. The clusters of AEs produced by our DP priors have not been investigated but may reveal additional, deeper insights by identifying groups of “similar AEs”. A reformulation via mixed membership models could further enhance interpretability by allowing the same AE to appear in multiple clusters/groups. Second, while we consider data-driven characterization of similarities in AE signal strengths between drugs, a more flexible approach acknowledging known Drug-Drug Interactions (DDI Poleksic and Xie (2019)), possibly via appropriate prior distributions, may further enhance signal detection. Finally, our current framework aggregates drug-AE data across time points for a cross-sectional analysis. Extending the model to accommodate temporal variations in SRS data could enable formal identification of evolving AE patterns.
@@
$\lambda_{ij}\sim z_{ij}\lambda_{ij}^{l}+(1-z_{ij})\lambda_{i}^{g}$
@@
$\lambda_{ij}\sim z_{ij}\lambda_{ij}^{l}+(1-z_{ij})\lambda_{i}^{g}$
@@
In this section, we first provide MCMC algorithms implementing our models in Section S.2.1. The MCMC algorithms are derived using finite mixture approximation of the underlying DP processes [23, 16], and combine Gibbs sampling updates with stepping-out slice sampling updates [31] for various parameters. Some additional notes on the MCMC implementation, including initial value generation for our MCMC algorithms are then provided in Section S.2.2. Next we provide detailed notes on the grid-search method identifying optimal quantile-level $p$ and threshold $k$ used for our hypothesis tests for signal detection (Section S.2.3). The JAGS code [33] for Hu et al. [18]’s models is shown as a reference in Section S.2.4. The example statin dataset is included in Section S.2.5.
@@
    Here $f_{\text{Gamma}}(\theta\mid\text{shape}=\beta,\text{rate}=\beta)$ denotes the probability density function of the Gamma($\beta,\beta$) distribution evaluated at $\theta$. $\pi(\beta)$ is the prior distribution of $\beta$. Direct random sampling from this density is not feasible. Instead, we use stepping-out slice sampling [31] to generate Markov chain draws for $\beta$.
@@
Algorithm S.2 One Step Iteration of DP Poisson local-global column model (Given data $\{(n_{ij},E_{ij},z_{ij}):i=1,\dots,I;j=1,\dots,J\}$ )
@@
    Generate global $\text{DPSB}_{K}(\bm{\eta},\bm{\theta},\alpha,\beta)$ parameters using Algorithm S.1, with $\{(\tilde{n}_{i},\tilde{E}_{i},1(\sum_{j=1}^{J-1}z_{ij}<J-1)):i=1,\dots,I\}$ as data where
    $\tilde{n}_{i}=\sum_{j=1}^{J-1}n_{ij}(1-z_{ij}),\quad\tilde{E}_{i}=\sum_{j=1}^{J-1}E_{ij}(1-z_{ij}).$
@@
    $z_{ij}\mid n_{ij},E_{ij},z_{ij},\lambda_{ij}^{l},\lambda_{i}^{g}\sim\text{Bernoulli}(p_{ij}),$
@@
    $p_{ij}=\frac{\pi f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})}{\pi f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})+(1-\pi)f_{\text{Poisson}}(n_{ij}\mid\lambda_{i}^{g}E_{ij})}.$
@@
    Define $\lambda_{ij}=z_{ij}\lambda_{ij}^{l}+(1-z_{ij})\lambda_{i}^{g}$ for all $i=1,\dots,I$ and $j=1,\dots,(J-1)$.
@@
    Generate $\pi$ using:
    $\pi\mid z_{ij}\sim\text{Beta}\left(a_{\pi}+\sum_{i=1}^{I}\sum_{j=1}^{J-1}z_{ij},b_{\pi}+\sum_{i=1}^{I}\sum_{j=1}^{J-1}(1-z_{ij})\right),$
    Notes: $a_{\pi}$ = 1, $b_{\pi}$ = 1.
@@
Algorithm S.3 One Iteration of DP ZIP local-global Model Given data $(n_{ij},E_{ij},z_{ij})$
@@
Given the observed data $n_{ij}$ and $E_{ij}$ for $i=1,\dots,I$, $j=1,\dots,J$, and a prespecified number $K$ of DP mixture components for finite approximation. The default value for $K$ in our software is set to $\max(\log(IJ),10)$, a setting that is generally sufficient for most cases. For the simulation study in Section 5, we used $K=10$, which satisfies the condition $K>\log(1492\times 7)$. In the data analysis presented in Section 6, we chose a more conservative value of $K=20$ to allow for a larger number of atoms in the DP, facilitating a more cautious interpretation of the results. We initialize the number of nonempty components $K_{+}$ to the smallest integer greater than $\min\{K/2,I-1\}$. The local-global indicators $z_{ij}$ are randomly initialized by independently sampling each cell $(i,j)$ from a $\text{Bernoulli}(0.5)$ distribution. The DP parameters for each local and global DP process are then initialized using the subset of data $(n_{ij},E_{ij})$ assigned to that process, as determined by the indicators $z_{ij}$, for $i=1,\dots,J-1$.
@@
Table S.1: The first 30 rows of the SRS contingency table summarized from the FDA FEARS database (Q3 2014-Q4 2020) which catalogs a total of 63,976,610 reported cases. There are 1490 preferred terms in the full dataset.
@@
We also present the full dataset with a smaller setting involving a subset of Table S.1 with only $I=46$ relevant AEs as considered in prior work[9, 19]. It is used for data generation in the simulation of Section S.3.4.
Table S.2: The SRS contingency table summarized from the FDA FEARS database (Q3 2014-Q4 2020) which catalogs a total of 63,976,610 reported cases. There are 46 preferred terms in the full dataset.
Acute Kidney Injury 1353 42 7 154 689 823 355651
Anuria 71 0 0 2 43 62 10403
Blood Calcium Decreased 14 2 0 0 110 17 15918
Blood Creatine Phosphokinase Abnormal 34 0 0 0 8 11 261
Blood Creatine Phosphokinase Increased 1175 125 32 200 562 768 23805
Blood Creatine Phosphokinase Mm Increased 2 0 0 0 9 0 14
Blood Creatinine Abnormal 27 0 0 0 5 3 3385
Blood Creatinine Increased 227 10 0 17 210 97 74742
Chromaturia 340 10 6 33 174 114 19294
Chronic Kidney Disease 152 16 2 19 177 37 339179
Compartment Syndrome 53 0 0 1 12 12 2644
Creatinine Renal Clearance Decreased 6 0 0 2 124 6 7768
Diaphragm Muscle Weakness 14 0 0 0 8 1 94
Electromyogram Abnormal 2 0 0 2 0 0 132
End Stage Renal Disease 30 0 0 0 19 6 97553
Glomerular Filtration Rate Abnormal 8 0 0 0 1 0 1069
Glomerular Filtration Rate Decreased 59 1 0 8 39 29 13190
Hypercreatininaemia 0 0 0 0 8 0 648
Hypocalcaemia 36 0 0 16 8 18 23102
Muscle Disorder 291 2 7 21 191 87 7329
Muscle Enzyme Increased 48 1 0 0 13 9 410
Muscle Fatigue 85 0 2 16 30 39 4257
Muscle Haemorrhage 24 0 0 5 13 4 3806
Muscle Necrosis 68 2 0 1 10 20 662
Muscle Rupture 181 25 0 61 36 120 3219
Muscular Weakness 1857 45 31 152 808 859 111003
Musculoskeletal Discomfort 137 18 15 25 187 93 19931
Musculoskeletal Disorder 56 3 0 9 65 73 25881
Musculoskeletal Pain 420 3 2 38 324 228 82576
Myalgia 5362 341 151 939 2757 3216 143819
Myasthenic Syndrome 1 0 9 0 3 6 643
Myoglobin Blood Increased 71 4 0 4 16 39 1003
Myoglobin Blood Present 2 0 0 0 0 0 8
Myoglobin Urine Present 0 0 0 0 1 2 70
Myoglobinaemia 15 0 0 0 0 0 62
Myoglobinuria 26 4 0 1 1 9 296
Myopathy 849 64 45 145 377 544 6695
Myopathy Toxic 31 0 0 1 2 21 457
Myositis 219 8 10 28 62 141 7482
Necrotising Myositis 279 0 0 2 10 52 278
Oliguria 52 0 0 4 24 37 7590
Renal Failure 534 26 11 69 225 195 250710
Renal Impairment 390 52 11 37 161 169 103343
Renal Tubular Necrosis 40 0 0 12 10 24 12762
Rhabdomyolysis 2041 52 44 163 936 1376 31707
Tendon Discomfort 9 0 0 3 22 10 794
Other Pt 180699 4886 2845 20296 113960 76068 61724222
@@
We present detailed results from our simulation experiments to compare among our local-global models (Poisson/ZIP), DP Hu et al. model, and a posterior quantile-based, FNR-optimized DP Hu et al. test method. Four evaluation metrics (FDR, sensitivity, average type I error, and F-score) are reported under the same simulation settings in Table 1 of the main test. Frequentist and Bayesian MSE are compared to further evaluate the loss of the estimation.
@@
We present the four evaluation metrics in Tables S.3, S.4, S.5, and S.6 to provide a clearer numerical comparison beyond Figure 3. Additionally, we further explored the effect of posterior quantile optimization on the DP Hu et al. in the process of FDR control and minimizing FNR. This allows for a more comprehensive understand the strategy of optimizing the posterior quantile of the test statistic and deterministic threshold simultaneously.
@@
Case 0a No correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
1.1 0.0000 0.0000 0.0000 0.0000
1.2 0.0136 0.0142 0.0047 0.0165
1.3 0.0152 0.0148 0.0088 0.0161
1.4 0.0119 0.0106 0.0049 0.0117
1.5 0.0088 0.0090 0.0046 0.0120
1.7 0.0097 0.0097 0.0044 0.0112
1.9 0.0066 0.0068 0.0037 0.0072
2 0.0073 0.0074 0.0043 0.0090
2.5 0.0060 0.0060 0.0032 0.0062
3 0.0048 0.0050 0.0023 0.0056
Case 0b No correlation 30% ZI 1.01 0.0050 0.0000 0.0050 0.0150
1.1 0.0052 0.0052 0.0103 0.0155
1.2 0.0105 0.0093 0.0044 0.0155
1.3 0.0101 0.0105 0.0077 0.0146
1.4 0.0102 0.0092 0.0056 0.0125
1.5 0.0113 0.0102 0.0053 0.0142
1.7 0.0079 0.0078 0.0043 0.0098
1.9 0.0075 0.0075 0.0033 0.0088
2 0.0070 0.0071 0.0032 0.0074
2.5 0.0068 0.0067 0.0035 0.0076
3 0.0056 0.0055 0.0025 0.0063
Case 1a Weak correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0155
@@
3 0.0042 0.0042 0.0026 0.0056
Case 1b Weak correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0100
@@
3 0.0037 0.0037 0.0022 0.0049
Case 2a Moderate correlation No ZI 1.01 0.0050 0.0050 0.0000 0.0000
@@
3 0.0024 0.0023 0.0026 0.0056
Case 2b Moderate correlation 30% ZI 1.01 0.0100 0.0050 0.0100 0.0100
@@
3 0.0025 0.0025 0.0029 0.0062
Case 3a Strong correlation No ZI 1.01 0.0051 0.0051 0.0000 0.0152
@@
3 0.0015 0.0017 0.0026 0.0058
Case 3b Strong correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0050
@@
Case 0a No correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
1.1 0.0005 0.0004 0.0001 0.0003
1.2 0.0269 0.0262 0.0202 0.0248
1.3 0.0734 0.0727 0.0582 0.0656
1.4 0.1199 0.1201 0.0985 0.1082
1.5 0.1605 0.1606 0.1366 0.1483
1.7 0.2269 0.2267 0.2012 0.2163
1.9 0.2848 0.2850 0.2596 0.2748
2 0.3098 0.3104 0.2849 0.3018
2.5 0.4169 0.4177 0.3939 0.4120
3 0.4921 0.4923 0.4711 0.4902
Case 0b No correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
1.1 0.0004 0.0004 0.0001 0.0003
1.2 0.0269 0.0273 0.0212 0.0252
1.3 0.0731 0.0732 0.0579 0.0662
1.4 0.1208 0.1203 0.0985 0.1099
1.5 0.1601 0.1601 0.1369 0.1492
1.7 0.2302 0.2297 0.2048 0.2192
1.9 0.2865 0.2868 0.2611 0.2769
2 0.3117 0.3115 0.2873 0.3038
2.5 0.4171 0.4164 0.3930 0.4115
3 0.4902 0.4903 0.4695 0.4875
Case 1a Weak correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
3 0.5245 0.5252 0.4818 0.4988
Case 1b Weak correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
3 0.5256 0.5261 0.4813 0.4998
Case 2a Moderate correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
3 0.6597 0.6599 0.5667 0.5826
Case 2b Moderate correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
3 0.6611 0.6613 0.5674 0.5835
Case 3a Strong correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
3 0.7959 0.7961 0.6307 0.6470
Case 3b Strong correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
2 0.0000 0.0000 0.0000 0.0000
2.5 0.0001 0.0001 0.0000 0.0001
3 0.0001 0.0001 0.0000 0.0001
Case 1a Weak correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
3 0.0000 0.0000 0.0000 0.0000
Case 1b Weak correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
1.4 0.0000 0.0000 0.0000 0.0000
1.5 0.0000 0.0000 0.0000 0.0000
1.7 0.0000 0.0000 0.0000 0.0000
1.9 0.0000 0.0000 0.0000 0.0000
2 0.0000 0.0000 0.0000 0.0000
2.5 0.0000 0.0000 0.0000 0.0000
3 0.0000 0.0000 0.0000 0.0000
Case 2b Moderate correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
1.1 0.0000 0.0000 0.0000 0.0000
1.2 0.0000 0.0000 0.0000 0.0000
1.3 0.0000 0.0000 0.0000 0.0000
1.4 0.0000 0.0000 0.0000 0.0000
1.5 0.0000 0.0000 0.0000 0.0000
1.7 0.0000 0.0000 0.0000 0.0000
1.9 0.0000 0.0000 0.0000 0.0000
2 0.0000 0.0000 0.0000 0.0000
2.5 0.0000 0.0000 0.0000 0.0000
3 0.0000 0.0000 0.0000 0.0000
Case 3a Strong correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
3 0.0000 0.0000 0.0000 0.0000
Case 3b Strong correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
Case 0a No correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
1.1 0.0010 0.0007 0.0002 0.0007
1.2 0.0520 0.0506 0.0394 0.0480
1.3 0.1359 0.1348 0.1094 0.1224
1.4 0.2132 0.2135 0.1787 0.1944
1.5 0.2755 0.2756 0.2395 0.2571
1.7 0.3686 0.3683 0.3340 0.3542
1.9 0.4419 0.4421 0.4111 0.4296
2 0.4715 0.4722 0.4423 0.4620
2.5 0.5868 0.5876 0.5640 0.5819
3 0.6580 0.6581 0.6394 0.6561
Case 0b No correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
1.1 0.0009 0.0008 0.0002 0.0007
1.2 0.0520 0.0527 0.0412 0.0488
1.3 0.1355 0.1356 0.1089 0.1235
1.4 0.2147 0.2139 0.1786 0.1971
1.5 0.2749 0.2749 0.2401 0.2584
1.7 0.3729 0.3722 0.3390 0.3582
1.9 0.4439 0.4443 0.4131 0.4321
2 0.4738 0.4735 0.4453 0.4645
2.5 0.5868 0.5861 0.5631 0.5811
3 0.6561 0.6562 0.6380 0.6535
Case 1a Weak correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
3 0.6864 0.6869 0.6489 0.6635
Case 1b Weak correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
3 0.6875 0.6879 0.6486 0.6646
Case 2a Moderate correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
3 0.7938 0.7939 0.7223 0.7342
Case 2b Moderate correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
3 0.7948 0.7949 0.7227 0.7347
Case 3a Strong correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
3 0.8856 0.8856 0.7725 0.7836
Case 3b Strong correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
@@
3 0.8857 0.8858 0.7718 0.7829
S.3.2 Simulation results for sensitivity analysis for grid configuration of $k$ in hypothesis testing
In our signal detection procedure, a sequence of threshold values for $k$ is constructed on a predefined grid. By default, we consider a grid ranging from $k=1.1$ to $k=3$, with small increments. This range is primarily motivated by the simulation design, where the maximum signal strength is set to be no larger than $3$, rendering this grid sufficient to capture all relevant signal levels under the data-generating mechanism.
To assess whether the default range for $k$ is restrictive, we conducted an additional sensitivity analysis under simulation Cases 2a. In this analysis, we adopted a highly conservative strategy by extending the upper bound of the grid to the maximum value observed among all MCMC draws of $\lambda$, effectively performing a global search over a substantially wider range of $k$. The resulting performance curves under this expanded grid are indistinguishable from those obtained using the default grid, across all evaluation metrics.
These results in Figure S.3.2 indicate that, for the considered simulation setting under Case 2a, the default grid range $(1.1,3)$ is adequate and does not lead to any loss in detection performance. More importantly, the consistency between the two strategies suggests that the proposed method is robust to the choice of the search range for $k$. Consequently, in applications involving datasets with potentially stronger signals, adopting a more conservative, data-driven upper bound for $k$ remains a safe option and does not adversely affect performance.
S.3.3 Simulation results for sensitivity analysis for hyperparameters
We conducted a simulation study under Case 2a to evaluate the robustness of the proposed DP Poisson local-global model with respect to the choice of hyperparameters. The model structure was kept fixed throughout the simulations, and sensitivity analysis was performed by varying the hyperparameters governing the prior distributions of $\alpha$ and $\beta$, which control the strength and concentration of shrinkage in the local-global framework.
Specifically, the baseline model adopts $\psi_{\alpha}=3$ and $\psi_{\beta}=0.5$ as our default model. To assess sensitivity, we considered alternative hyperprior specifications that represent substantially different prior beliefs while remaining within reasonable ranges. For $\alpha$, we examined a more informative prior with $\psi_{\alpha}=5$, placing approximately $99\%$ prior mass within $(0,1)$, as well as a weaker prior with $\psi_{\alpha}=1$, placing approximately $64\%$ mass in $(0,1)$. For $\beta$, we considered $\psi_{\beta}=1/7$, corresponding to a prior favoring values in $(0,1)$ with approximately $90\%$ prior mass in this interval, and $\psi_{\beta}=6$, which places only $10\%$ prior mass in $(0,1)$ and thus favors more extreme shrinkage behavior.
These hyperprior configurations are summarized in Table S.7. The comparison across these settings allows us to evaluate the stability of model performance under varying degrees of prior informativeness and shrinkage strength.
Table S.7: Hyperprior configurations used in the sensitivity analysis. The baseline setting corresponds to the default hyperprior choice in the proposed DP Poisson local–global model.
Model setting $\psi_{\alpha}$ $\psi_{\beta}$ Interpretation
Baseline 3 0.5 Default hyperprior specification in the proposed model
$\alpha$-strong 5 0.5 More informative prior on $\alpha$, placing $\approx 99\%$ mass in $(0,1)$
$\alpha$-weak 1 0.5 Weaker prior on $\alpha$, placing $\approx 64\%$ mass in $(0,1)$
$\beta$-favoring 3 $1/7$ Prior favoring $\beta\in(0,1)$ with $\approx 90\%$ mass in $(0,1)$
$\beta$-sparse 3 6 Prior discouraging $\beta\in(0,1)$, placing only $10\%$ mass in $(0,1)$
Figure S.1: Simulation results under different hyperprior configurations for the DP Poisson local-global model. Curves corresponding to all hyperprior settings largely overlap across performance metrics, indicating robust model performance with respect to reasonable variations in hyperparameter choices.
As shown in Figure S.1, across all performance metrics, the curves corresponding to different hyperprior configurations largely overlap in the simulation results. This indicates that the proposed DP Poisson local-global model exhibits stable performance under a wide range of hyperparameter choices. In particular, varying the hyperpriors for $\alpha$ and $\beta$, from relatively informative to weakly informative specifications, does not materially affect model behavior at the considered data scale. These findings suggest that the proposed model is reasonably robust to reasonable variations in hyperprior assumptions. Moreover, the sensitivity analysis covers several representative hyperprior scenarios, which together are expected to reflect a broad range of practical data-generating settings.
S.3.4 Simulation results under smaller setting
We conduct the simulation based on Table S.2 under similar correlation strength setting in Table 1. The result shown in Figure S.2 has the same pattern as we observed in Section 5 based on larger setting.
Figure S.2: Simulation results: the horizontal axis shows the true signal strength $\lambda_{0}$ for signal cells, the vertical axis shows the value of evaluation metrics. Each column presents one metric and each row presents one simulation scenario.